OpenStudy (anonymous):

I need an explanation on how to do problem like this Find the first 5 terms of the sequence: an = 4^n - 1.

OpenStudy (across):

A sequence is usually defined over the set of natural numbers \(\mathbb{N}\). It follows that all you have to do is know how to plug in numbers: Given\[a_n=4^n-1,\]we can see that the first \(5\) terms of this function are \(3,15,63,255,1023\).

OpenStudy (across):

I meant to say sequence, not function. :P

OpenStudy (anonymous):

so I would plug in the real numbers for n and then do the math?

OpenStudy (across):

The natural numbers. :)

OpenStudy (across):

Recall that the natural numbers are\[\mathbb{N}=\{1,2,3,\dots\}.\]

OpenStudy (anonymous):

oh yeah.. oops lol

OpenStudy (anonymous):

okay, thats easy enough. Thanks(: